OpenStudy (anonymous):

Determine # of solutions: 1/3x2- 5x +25

OpenStudy (anonymous):

Sorry =0*

jimthompson5910 (jim_thompson5910):

From (1/3)x^2-5x+25, we see that a = 1/3, b = -5 and c = 25 D=b^2-4ac D=(-5)^2-4(1/3)(25) D=25-4(1/3)(25) D=25-100/3 D=-25/3 So the discriminant is D=-25/3 Since the discriminant is less than zero, this means that there are two complex solutions. In other words, there are no real solutions.

OpenStudy (anonymous):

So the formula to find a discriminant is b^2-4ac?

jimthompson5910 (jim_thompson5910):

yep, that's it

OpenStudy (anonymous):

Thanks so much!